OpenStudy (anonymous):

Which of the following equations is of a parabola with a vertex at (0, 3)?

OpenStudy (anonymous):

k what are the equations? ().()?

OpenStudy (anonymous):

?.?

OpenStudy (anonymous):

you can copy and paste the quistion

OpenStudy (anonymous):

a. y=(x-3)^2 b. y=(x+3)^2 c. Y=x^2-3 d. y=x^2+3